• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Xiaohui

    Log entry time 00:02:50 on May19,2010

    Entry number 323194

    keyword=Shift Summary

    Shift summary:

    16:00 Shift starts. Waiting for the beam. Septum is ON, target in empty.

    16:25 MCC almost ready to deliver the beam. Ask for 5uA, prepare for beam recovery.

    17:00 Switch to thick carbon target, and ask for 50 uA to heat the target. Beam position is good.

    17:48 Moved to empty target.

    17:50 Back to beam recovery, target moved to BeO.

    20:12 Almost finish the beam recovery, target moved to Pb#2, ask for 5uA.

    20:35 Ramped the current to 50uA, start production.

    production runs:
    run# 3936 left Q1 tripped in the end of the run
    run# 3937
    run# 3938
    22:34 Stop production, Rupesh starts to do the over-sampleing test.